On January 1, 2021, Splash City issues $470,000 of 9% bonds, due in 20 years, with interest payable semiannually on June 30 and December 31 each year. Assuming the market interest rate on the issue date is 10%, the bonds will issue at $429,678. Required: 1. Complete the first three rows of an amortization table.
Answer #1

Par value of bonds = $470,000

Semi annual interest payment = 470,000 x 9% x 6/12

= $21,150

Effective interest rate = 10%

Semi annual Effective interest rate = 5%

Issue price of bonds = $429,678

Discount on bonds payable = Par value of bonds - Issue price of bonds

= 470,000 - 429,678

= $40,322

Date

Cash paid

Interest expense

Change in carrying value

Carrying value

1/1/21

429,678

6/30/21

21,150

429,678 x 5% = 21,484

21,484 - 21,150 = 334

429,678 + 334 = 430,012

12/31/21

21,150

430,012 x 5% = 21,501

21,501 - 21,150 = 351

430,012 + 351 = 430,363

Please ask if you have any query related to the question. Thank you